# Density estimation with LLMs: a geometric investigation of in-context learning trajectories

### v2

@@ Abstract
Large language models (LLMs) demonstrate remarkable emergent abilities to perform in-context learning across various tasks, including time series forecasting. This work investigates LLMs’ ability to estimate probability density functions (PDFs) from data observed in-context; such density estimation (DE) is a fundamental task underlying many probabilistic modeling problems. We leverage the Intensive Principal Component Analysis (InPCA) to visualize and analyze the in-context learning dynamics of LLaMA-2 models. Our main finding is that these LLMs all follow similar learning trajectories in a low-dimensional InPCA space, which are distinct from those of traditional density estimation methods like histograms and Gaussian kernel density estimation (KDE). We interpret the LLaMA in-context DE process as a KDE with an adaptive kernel width and shape. This custom kernel model captures a significant portion of LLaMA’s behavior despite having only two parameters. We further speculate on why LLaMA’s kernel width and shape differs from classical algorithms, providing insights into the mechanism of in-context probabilistic reasoning in LLMs.
@@ 1 Introduction
Recently, LLMs have been shown to achieve competitive performance on various mathematical problems, including time-series forecasting (Gruver et al., 2024), inferring physical rules from dynamical systems (Kantamneni et al., 2024; Liu et al., 2024), and learning random languages (Bigelow et al., 2024). To solve these tasks, an LLM must possess some capacity for probabilistic modeling — the ability to infer conditional or unconditional probability distribution structures by collecting relevant statistics from in-context examples (Akyürek et al., 2024).
We investigate LLMs’ ability to perform density estimation (DE), which involves estimating the probability density function (PDF) from data observed in-context. Our core experiment is remarkably straightforward. As illustrated in Figure 1, we prompt LLaMA-2 models (Touvron et al., 2023) with a series of data points $\{X_{i}\}_{i=1}^{n}$ sampled independently and identically from an underlying distribution $p(x)$. We then observe that the LLaMA’s predicted PDF, $\hat{p}_{n}(x)$, for the next data point gradually converges to the ground truth as the context length $n$ (the number of in-context data points) increases.. [^1] Our prompts consist purely of comma-delimited multi-digit numbers, without any natural language instructions.
To interpret the internal mechanisms (Olsson et al., 2022; Bietti et al., 2023; Dai et al., 2023; von Oswald et al., 2023) underlying an LLM’s in-context DE process, we leverage Intensive Principal Component Analysis (InPCA) (Teoh et al., 2020; Quinn et al., 2019; Mao et al., 2024) to embed the estimated PDF at each context length $\{\hat{P}_{n}\}$ in probability space (Figure 2). These visualizations reveal that the in-context DE process of these models all follow similar low-dimensional paths, which are distinct from those of traditional density estimation methods like histograms and Gaussian kernel density estimation (KDE).
By studying the geometric features of these in-context DE trajectories, we identify a strong bias towards Gaussianity, which we argue is a telltale feature of kernel-based density estimationRosenblat (1956); Wand & Jones (1994); Silverman (2018). This observation inspires us to model LLaMA in-context DE process as a kernel density estimator with adaptive kernel. Despite having only two parameters, kernel shape and width, this bespoke KDE model captures LLaMA’s in-context learning trajectories with high precision.
Figure 1: In-context density estimation experiment. LLaMA-2 is prompted with 200 numbers sampled from $p(x)$ (left in red) (Appendix A.7), and predicts the PDF $\hat{p}(x)$ (right in blue) for the next number.
@@
    We apply InPCA to analyze the low-dimensional geometric features of LLaMA-2’s in-context density estimation trajectories.
@@
    We interpret LLaMA-2’s in-context DE algorithm as a form of adaptive kernel density estimation, providing evidence for a dispersive induction head mechanism (Olsson et al., 2022; Akyürek et al., 2024).
@@ 2 Background
Gruver et al. (2024) pioneered LLM-based time-series prediction by prompting LLMs with time-series data serialized as comma-delimited, multi-digit numbers. They observed competitive predictions for future states encoded as softmax probabilities. Their data serialization scheme proved simple yet effective, allowing for easy log-likelihood evaluation of predicted next numbers. It has since become a common prompting method (Jin et al., 2024; Requeima et al., 2024; Zhang et al., 2024; Liu et al., 2024) for evaluating LLMs’ numerical abilities.
@@
DE (Izenman, 1991) is a long-standing statistical problem with both classical (Silverman, 2018; Rosenblat, 1956; Scott, 1979; Lugosi & Nobel, 1996; Parzen, 1962) and modern machine-learned based solutions Papamakarios et al. (2021); Sohl-Dickstein et al. (2015). It underlies the learning of more complex stochastic processes. For example, in the first-order Markov process studied in (Liu et al., 2024; Bigelow et al., 2024; Zekri et al., 2024), the transition probability $p(x_{t}|x_{t-1})$ has to be estimated in-context. This can be viewed as a conditional density estimation problem, where for each possible value of $x_{t-1}$, one needs to estimate the density of $x_{t}$. In other words, learning a Markov process involves performing multiple density estimations; one for each conditioning state. Our current focus on unconditional density estimation thus serves as a stepping stone towards understanding these more complex learning tasks.
Figure 2: Visualization pipeline of LLaMA’s in-context density estimation process. (a) Data points are independently sampled from a ground truth distribution (Gaussian in this example), then serialized as comma-delimited, two-digit numbers to prompt LLaMA-2 models. (b) Hierarchy-PDF extracts LLaMA’s estimated density function $\hat{p}_{n}$ at each context length $n$. (c) InPCA reveals low-dimensional structures in density estimation trajectories, capturing 91% of pairwise Hellinger distances in 2 dimensions. Visual guides: gray - uniform PDF representing maximal ignorance, deep blue - ground truth PDF, and pink - 1D submanifold of centered Gaussians with variances ranging from $\infty$ to 0. All three LLaMA-2 models exhibit similar DE trajectories, geometrically bounded between the geodesic and the Gaussian sub-manifold.
@@ 3 Methodology
Here, we explain how we prompt a LLaMA-2 model to perform density estimation over in-context data, and how we extract and analyze its in-context learning trajectories with InPCA. Our methodology consists of 5 steps, the first 3 of which are illustrated in Fig. 2.
@@
2. Extracting estimated densities with Hierarchy-PDF. Upon prompting with such a text string, we read out LLaMA’s softmax prediction over the next token, yielding probabilities for 10 tokens (0-9) [^2] The logits of all other tokens are ignored., creating a coarse, 10-bin PDF spanning $x\in(0,100)$. We then read out the next token, which refines one of the tens bins by further dividing it into 10 ones bins. This process is repeated recursively 10 times, until all bins are refined, yielding a predicted PDF for the next state $p(x_{t})$, which is a discrete PDF object consisting of $10^{N}$ bins, where $N$ is the number of digits used in representing each number. [^3] We use the Hierarchy-PDF algorithm (Liu et al., 2024), which performs this recursive search efficiently for transformers. We interpret the extracted PDF $\hat{p}(x_{t})$ as the LLM’s estimation of the ground-truth $p(x)$.
3. Visualizing DE trajectories with InPCA. When there are very few in-context data, LLaMA’s estimated $\hat{p}_{0}(x)$ is close to a uniform distribution over the domain $(0,100)$, showing a state of neutral ignorance, which is a reasonable Bayesian prior (Xie et al., 2022). However, as the number of in-context data ($n$) increases, LLaMA gathers more information about the underlying distribution. As a result, the estimated $\hat{p}_{n}(x)$ gradually converges to the ground truth $p(x)$ (Figure 2 (b)). The series of estimated PDFs $\hat{p}_{0}(x),\hat{p}_{1}(x),...,\hat{p}_{n}(x)$ over context length $n$ forms what we term the "in-context DE trajectory". These trajectories, we argue, offer important clues about how LLaMA-2 performs in-context learning.
@@
4. Comparing with classical DE algorithms. To interpret the low-dimensional geometric features revealed by InPCA, we embed the DE trajectories of well-known algorithms — specifically, kernel density estimator (KDE) and Bayesian histograms — in the same two-dimensional (2D) space. Surprisingly, we find that the DE trajectories of LLaMA-2, KDE, and Bayesian histograms are simultaneously embeddable in this same low-dimensional subspace. Moreover, the DE trajectories of LLaMA-2 are geometrically bounded between those of KDE and Bayesian histograms.
5. Explaining LLaMA-2 with bespoke KDE. We observe that LLaMA’s in-context DE trajectories consistently remain close to those of KDE, suggesting that LLaMA might be employing a kernel-like algorithm, internally. To investigate this hypothesis, we optimize a KDE algorithm with fitted kernel parameters (using only two parameters).
@@ 3.1 InPCA visualization
The Hellinger distance is ideal for visualizing our PDF trajectories for the following reasons [^6] See Section A.5.3 for an in-depth comparison of the Hellinger distance against other divergence measures.: 1) it locally agrees with the KL-divergence (Liese & Vajda, 2006), a non-negative measure commonly used in training modern machine learning systems, including LLMs (Touvron et al., 2023), and 2) it is symmetric and satisfies the triangle inequality, making it a proper distance metric suitable for geometric analysis (Quinn et al., 2019; Teoh et al., 2020).
@@ 3.2 Classical DE algorithms
where $n$ is the sample size and $C$ is a pre-coefficient (Equation 13). In the low-data regime, a larger bandwidth provides more smoothing bias, which compensates for data sparsity. While the $n^{-1/5}$ scaling is widely accepted, determining the pre-coefficient $C$ is more challenging (Wand & Jones, 1994; Silverman, 2018). Unless otherwise noted, we use $C=1$ for classical KDE in this paper [^7] In practice, various heuristics have been proposed to determine $C$, such as Silverman’s rule of thumb (Silverman, 2018) (see Appendices A.5.1 and A.8)..
@@ 4 Experiments and analysis
We visualize and analyze the learning trajectories of LLaMA on two types of target (ground truth) distributions: uniform and Gaussian. To provide context and facilitate interpretation of the 2D space, we embed the following additional reference points and trajectories:
@@
  • Truth: A point representing the ground truth distribution.
  • Geodesic: The shortest trajectory connecting Ignorance and Truth points (Mao et al., 2024).
@@
Our experiments focus on LLaMA-2 13b, unless otherwise noted. While LLaMA-2 has a context window of 4096 tokens (equivalent to $\sim$1365 comma-delimited, 2-digit data points), we limit our analysis to a context length of $n=200$. This limitation is based on our observation that LLaMA’s DE accuracy typically plateaus beyond this point.
@@ 4.2 Uniform target
Narrow uniform target. As the uniform target narrows, Gaussian KDE’s performance deteriorates, while LLaMA’s in-context DE successfully reaches the target. Notably, LLaMA exhibits less bias towards the Gaussian sub-manifold, as compared with the narrow Gaussian target case.
In the low data regime, Gaussian KDE and Bayesian histogram follow the same trajectories that they previously followed for the narrow Gaussian target. However, LLaMA already appears to differentiate this target by taking a path further from the Gaussian sub-manifold and closer to the geodesic. This behavior suggests that LLaMA’s in-context DE algorithm is more flexible and adaptive than classical KDE with pre-determined width schedule and shape.
@@ 4.3.1 Bespoke KDE Model
yielding the "bespoke KDE" bandwidth schedule $\{h_{i}\}_{i=1}^{n}$ and shape schedule $\{s_{i}\}_{i=1}^{n}$, which together prescribe a sequence of fitted kernels of changing widths and shapes. We visualize such sequences of fitted kernels, and compare them against the Gaussian kernel with standard $n^{-1/5}$ width schedule in Appendix A.6 and A.7. We implement this optimization numerically using SciPy, and estimate parameter uncertainties from the inverse Hessian of the loss function (Hellinger distance) at the optimum. This provides error bars for our fitted kernel shape and bandwidth parameters (Figure 7).
@@ 4.3.2 Bespoke KDE Analysis
Significance of Bespoke KDE The existence of a bespoke KDE that closely imitates LLaMA’s in-context DE processs is significant and non-trivial. Not all DE processes can be faithfully represented by KDE methods; for instance, the Bayesian histogram, despite its superficial similarity to a fixed-width tophat kernel KDE, resists accurate modeling as a bespoke KDE process (Appendix A.3).
While our initial experiments focused on regular distributions like Gaussian and uniform, for better generality we extended our experiments to randomly generated PDFs (Appendix A.7). Notably, LLaMA-13b’s DE performance on these irregular distributions can still be effectively described using our bespoke KDE process (Appendix A.6 and A.7). Although the DE trajectories of these random PDFs lack low-dimensional embeddings, and thus provide less geometric insight, they nonetheless corroborate our findings from the cases with more regular distributions.
This consistency across diverse target PDF types strengthens our core hypothesis about LLaMA’s underlying DE mechanism. Specifically, we posit that LLaMA’s in-context DE algorithm shares fundamental characteristics with kernel density estimation, albeit with adaptive kernel shape and bandwidth schedules that distinguish it from classical KDE approaches.

### v3

@@ Abstract
Large language models (LLMs) demonstrate remarkable emergent abilities to perform in-context learning across various tasks, including time-series forecasting. This work investigates LLMs’ ability to estimate probability density functions (PDFs) from data observed in-context; such density estimation (DE) is a fundamental task underlying many probabilistic modeling problems. We use Intensive Principal Component Analysis (InPCA) to visualize and analyze the in-context learning dynamics of LLaMA-2, Gemma, and Mistral. Our main finding is that these LLMs all follow similar learning trajectories in a low-dimensional InPCA space, which are distinct from those of traditional density estimation methods such as histograms and Gaussian kernel density estimation (KDE). We interpret these LLMs’ in-context DE process as a KDE algorithm with adaptive kernel width and shape. This custom kernel model captures a significant portion of the LLMs’ behavior despite having only two parameters. We further speculate on why the LLMs’ kernel width and shape differ from classical algorithms, providing insights into the mechanism of in-context probabilistic reasoning in LLMs. Our codebase, along with a 3D visualization of an LLM’s in-context learning trajectory, is publicly available at https://github.com/AntonioLiu97/LLMICL_inPCA.
@@ 1 Introduction
Recently, LLMs have been shown to achieve competitive performance in various mathematical problems, including time series forecasting (Gruver et al., 2024), inferring physical rules from dynamical systems (Kantamneni et al., 2024; Liu et al., 2024), and learning random languages (Bigelow et al., 2024). To solve these tasks, an LLM must possess some capacity for probabilistic modeling — the ability to infer conditional or unconditional probability distribution structures by collecting relevant statistics from in-context examples (Akyürek et al., 2024).
We investigate LLMs’ ability to perform density estimation (DE), which involves estimating the probability density function (PDF) from data observed in-context. Our core experiment is remarkably straightforward. As illustrated in Figure 1, we prompt LLMs such as LLaMA-2 (Touvron et al., 2023), Gemma (GemmaTeam et al., 2024), and Mistral (Jiang et al., 2023) with a series of data points $\{X_{i}\}_{i=1}^{n}$ sampled independently and identically from an underlying distribution $p(x)$. We then observe that the LLMs’ predicted PDF, $\hat{p}_{n}(x)$, for the next data point gradually converges to the ground truth as the context length $n$ (the number of in-context data points) increases. [^1] Our prompts consist purely of comma-delimited multi-digit numbers, without any natural language instructions.
To interpret the internal mechanisms (Olsson et al., 2022; Bietti et al., 2023; Dai et al., 2023; von Oswald et al., 2023) underlying an LLM’s in-context DE process, we use Intensive Principal Component Analysis (InPCA) (Teoh et al., 2020; Quinn et al., 2019; Mao et al., 2024) to embed the estimated PDF at each context length $\{\hat{P}_{n}\}$ in the probability space (Figure 2). Our visualizations reveal that the in-context DE processes of the tested LLMs all follow similar low-dimensional paths, which are distinct from those of traditional density estimation methods like histograms and Gaussian kernel density estimation (KDE).
By studying the geometric features of these in-context DE trajectories, we identify a strong bias towards Gaussianity, which we argue is a telltale feature of kernel-based density estimation (Rosenblat, 1956; Wand & Jones, 1994; Silverman, 2018). This observation inspires us to model an LLM’s in-context DE process as a kernel density estimator with adaptive kernel. Despite having only two parameters, kernel shape and width, this bespoke KDE model captures the in-context learning trajectories with high precision.
Figure 1: In-context density estimation experiment. LLaMA-2 13b is prompted with 200 numbers sampled from $p(x)$ (left in red) (Appendix A.9), and predicts the PDF $\hat{p}(x)$ (right in blue) for the next number.
@@
    We apply InPCA to analyze the low-dimensional features in the in-context learning trajectories of various LLMs.
@@
    We interpret the LLMs’ in-context DE algorithm as a form of adaptive kernel density estimation, providing evidence for a dispersive induction head mechanism (Olsson et al., 2022; Akyürek et al., 2024).
@@ 2 Background
Gruver et al. (2024) pioneered LLM-based time-series prediction by prompting an LLM with time-series data serialized as comma-delimited, multi-digit numbers. They observed competitive predictions for future states encoded as softmax probabilities. Their data serialization scheme proved simple yet effective, allowing for easy log-likelihood evaluation of predicted next numbers. It has since become a common prompting method (Jin et al., 2024; Requeima et al., 2024; Zhang et al., 2024; Liu et al., 2024) for evaluating LLMs’ numerical abilities.
@@
DE (Izenman, 1991) is a long-standing statistical problem with both classical (Silverman, 2018; Rosenblat, 1956; Scott, 1979; Lugosi & Nobel, 1996; Parzen, 1962) and modern machine-learned solutions (Papamakarios et al., 2021; Sohl-Dickstein et al., 2015). It underlies the learning of more complex stochastic processes. For example, in the first-order Markov process studied in (Liu et al., 2024; Bigelow et al., 2024; Zekri et al., 2024), the transition probability $p(x_{t}|x_{t-1})$ has to be estimated in context. This can be viewed as a conditional density estimation problem, where for each possible value of $x_{t-1}$, the density of $x_{t}$ must be estimated. In other words, learning a Markov process involves performing multiple density estimations; one for each conditioning state. Our current focus on unconditional density estimation thus serves as a stepping stone towards understanding these more complex learning tasks.
Figure 2: Visualization pipeline of LLMs’ in-context density estimation process. (a) Data points are independently sampled from a ground truth distribution (Gaussian in this example), then serialized as comma-delimited, two-digit numbers to prompt LLMs, such as LLaMA-2, Mistral-v0.3, and Gemma. (b) Hierarchy-PDF extracts each LLM’s estimated density function $\hat{p}_{n}$ at each context length $n$. (c) InPCA reveals low-dimensional structures in density estimation trajectories, capturing 91% of pairwise Hellinger distances in 2 dimensions. Visual guides: gray - uniform PDF representing maximal ignorance, deep blue - ground truth PDF, and pink - 1D submanifold of centered Gaussians with variances ranging from $\infty$ to 0. All LLMs investigated in this work exhibit similar DE trajectories, which are geometrically bounded between the geodesic and the Gaussian sub-manifold.
@@ 3 Methodology
Here, we explain how we prompt an LLM to perform density estimation over in-context data, and how we extract and analyze its in-context learning trajectories with InPCA. Our methodology consists of 5 steps, the first 3 of which are illustrated in Fig. 2.
@@
2. Extracting estimated densities with Hierarchy-PDF. Upon prompting with such a text string, we read out the LLM’s softmax prediction over the next token, yielding probabilities for 10 tokens (0-9) [^2] The logits of all other tokens are ignored., creating a coarse, 10-bin PDF spanning $x\in(0,100)$. We then read out the next token, which refines one of the tens bins by further dividing it into 10 ones bins. This process is repeated recursively 10 times, until all bins are refined, yielding a predicted PDF for the next state $p(x_{t})$, which is a discrete PDF object consisting of $10^{N}$ bins, where $N$ is the number of digits used in representing each number. [^3] We use the Hierarchy-PDF algorithm (Liu et al., 2024), which performs this recursive search efficiently for transformers. We interpret the extracted PDF $\hat{p}(x_{t})$ as the LLM’s estimation of the ground-truth $p(x)$.
3. Visualizing DE trajectories with InPCA. When there are very few in-context data, an LLM’s estimated $\hat{p}_{0}(x)$ is close to a uniform distribution over the domain $(0,100)$, showing a state of neutral ignorance, which is a reasonable Bayesian prior (Xie et al., 2022). However, as the number of in-context data ($n$) increases, the LLM gathers more information about the underlying distribution. As a result, the estimated $\hat{p}_{n}(x)$ gradually converges to the ground truth $p(x)$ (Figure 2 (b)). The series of estimated PDFs $\hat{p}_{0}(x),\hat{p}_{1}(x),...,\hat{p}_{n}(x)$ over context length $n$ forms what we term the "in-context DE trajectory". These trajectories, we argue, offer important clues about how the LLM performs in-context learning.
@@
4. Comparing with classical DE algorithms. To interpret the low-dimensional geometric features revealed by InPCA, we embed the DE trajectories of well-known algorithms — specifically, kernel density estimator (KDE) and Bayesian histograms — in the same two-dimensional (2D) space. Surprisingly, we find that the DE trajectories of LLMs, KDE, and Bayesian histograms are simultaneously embeddable in this same low-dimensional subspace. Moreover, the DE trajectories of LLMs are geometrically bounded between those of KDE and Bayesian histograms.
5. Explaining LLaMA-2 with bespoke KDE. We observe that LLMs’ in-context DE trajectories consistently remain close to those of KDE, suggesting that LLMs might be employing a kernel-like algorithm, internally. To investigate this hypothesis, we optimize a KDE algorithm with fitted kernel parameters (using only two parameters).
@@ 3.1 InPCA visualization
The Hellinger distance is ideal for visualizing our PDF trajectories for the following reasons [^6] See Sections A.5 and A.6 for viusalizations using other distance measures such as L2 and the symmetrized KL-divergence.: 1) it locally agrees with the KL-divergence (Liese & Vajda, 2006), a non-negative measure commonly used in training modern machine learning systems, including LLMs (Touvron et al., 2023), and 2) it is symmetric and satisfies the triangle inequality, making it a proper distance metric suitable for geometric analysis (Quinn et al., 2019; Teoh et al., 2020).
@@ 3.2 Classical DE algorithms
where $n$ is the sample size and $C$ is a pre-coefficient (Equation 16). In the low-data regime, a larger bandwidth provides more smoothing bias, which compensates for data sparsity. While the $n^{-1/5}$ scaling is widely accepted, determining the pre-coefficient $C$ is more challenging (Wand & Jones, 1994; Silverman, 2018). Unless otherwise noted, we use $C=1$ for classical KDE in this paper [^7] In practice, various heuristics have been proposed to determine $C$, such as Silverman’s rule of thumb (Silverman, 2018) (see Appendices A.7.1 and A.13)..
@@ 4 Experiments and analysis
We visualize and analyze the learning trajectories of an LLM on two types of target (ground truth) distributions: uniform and Gaussian. To provide context and facilitate interpretation of the 2D space, we embed the following additional reference points and trajectories:
@@
  • Truth: A point representing the ground-truth distribution.
  • Geodesic: The shortest trajectory connecting the Ignorance and Truth points (Mao et al., 2024).
@@
For brevity, this section focuses on LLaMA-2 13B, as other LLMs yield qualitatively similar results, detailed in Appendix A.4. While LLaMA-2 has a context window of 4096 tokens (equivalent to $\sim$1365 comma-delimited, 2-digit data points), we limit our analysis to a context length of $n=200$. This limitation is based on our observation that LLaMA’s DE accuracy typically plateaus beyond this point.
@@ 4.2 Uniform target
Narrow uniform target. As the uniform target narrows, Gaussian KDE’s performance deteriorates, while LLM-based in-context DE successfully reaches the target. Notably, the LLM exhibits less bias towards the Gaussian sub-manifold, as compared with the narrow Gaussian target case.
In the low-data regime, Gaussian KDE and Bayesian histogram follow the same trajectories that they previously followed for the narrow Gaussian target. However, the LLM already appears to differentiate this target by taking a path further from the Gaussian sub-manifold and closer to the geodesic. This behavior suggests that LLMs’ in-context DE algorithm is more flexible and adaptive than classical KDE with pre-determined width schedule and shape.
@@ 4.3.1 Bespoke KDE Model
yielding the "bespoke KDE" bandwidth schedule $\{h_{i}\}_{i=1}^{n}$ and shape schedule $\{s_{i}\}_{i=1}^{n}$, which together prescribe a sequence of fitted kernels of changing widths and shapes. We visualize such sequences of fitted kernels, and compare them against the Gaussian kernel with standard $n^{-1/5}$ width schedule in Appendix A.8 and A.9. We implement this optimization numerically using SciPy, and estimate parameter uncertainties from the inverse Hessian of the loss function (Kl-divergence) at the optimum. This provides error bars for our fitted kernel shape and bandwidth parameters (Figure 7). In Section A.12, we elaborate the information-theoretic motivation for our uncertainty quantification method.
@@ 4.3.2 Bespoke KDE Analysis
Significance of Bespoke KDE The existence of a bespoke KDE that closely imitates an LLM’s in-context DE processs is significant and non-trivial. Not all DE processes can be faithfully represented by KDE methods; for instance, the Bayesian histogram, despite its superficial similarity to a fixed-width tophat kernel KDE, resists accurate modeling as a bespoke KDE process (Appendix A.3).
While our initial experiments focused on regular distributions like Gaussian and uniform, for better generality we extended our experiments to randomly generated PDFs (Appendix A.9). Notably, the LLMs’ DE performance on these irregular distributions can still be effectively described using our bespoke KDE process (Appendix A.8 and A.9). Although the DE trajectories of these random PDFs lack low-dimensional embeddings, and thus provide less geometric insight, they nonetheless corroborate our findings from the cases with more regular distributions.
This consistency across diverse target PDF types strengthens our core hypothesis about the in-context DE mechanism underlying LLMs. Specifically, we posit that an LLM’s in-context DE algorithm shares fundamental characteristics with kernel density estimation, albeit with adaptive kernel shape and bandwidth schedules that distinguish it from classical KDE approaches.